### v3

# CONQURE: A Co-Execution Environment for Quantum and Classical Resources

Abstract
Cutting edge classical computing today relies on a combination of CPU-based computing with a strong reliance on accelerators. In particular, high-performance computing (HPC) and machine learning (ML) rely heavily on acceleration via GPUs for numerical kernels. In the future, acceleration via quantum devices may complement GPUs for kernels where algorithms provide quantum advantage, i.e., significant speedups over classical algorithms. Computing with quantum kernels mapped onto quantum processing units (QPUs) requires seamless integration into HPC and ML. However, quantum offloading onto HPC/cloud lacks open-source software infrastructure. For classical algorithms, parallelization standards, such as OpenMP, MPI, or CUDA exist. In contrast, a lack of quantum abstractions currently limits the adoption of quantum acceleration in practical applications creating a gap between quantum algorithm development and practical HPC integration. Such integration needs to extend to efficient quantum offloading of kernels, which further requires scheduling of quantum resources, control of QPU kernel execution, tracking of QPU results, providing results to classical calling contexts and coordination with HPC scheduling.
This work proposes CONQURE, a co-execution environment for quantum and classical resources. CONQURE is a fully open-source cloud queue framework that presents a novel modular scheduling framework allowing users to offload OpenMP quantum kernels to QPUs as quantum circuits, to relay results back to calling contexts in classical computing, and to schedule quantum resources via our CONQURE API.
[…]
I Introduction
Quantum computing has the potential to disrupt classical computing for a number of fields. These fields include clinical research [1], optimization [2], high-energy physics [3], finance [4] and logistics [5]. The catalog of problems where quantum computing promises potential for significant speedup over its classical counterparts is only increasing. As such, quantum computing has reached an early (yet still experimental) state of maturity. A number of device technologies ranging from superconducting devices (e.g., IBM Q, Rigetti, QCI, OQC, Google, Amazon) to ion traps (IonQ, Quantinuum) to neutral atoms (QuEra, Pasqal) have become available through cloud access, e.g., via Amazon Braket [6], Azure Quantum [7], qBraid [8] or directly from the vendors. Existing commercial solutions feature quantum computing ecosystems on the basis of python packages to facilitate software development over a number of hardware devices. Today’s leading quantum computing hardware technologies include Superconducting qubits (IBM, Google, Amazon, Alice & Bob), Ion-Traps (IONQ, Quantinuum) and Neutral Atoms (QuEra Computing Inc.). Software libraries like Qiskit [9], Cirq [10], Tket [11], Pennylane [12] and CUDA-Q [13] enable users to design quantum workloads while lower-level control of the hardware is enabled by interfacing through proprietary software layers, or research infrastructure such as DAX [14], ARTIQ [15], Qubic [16, 17], and QICK [18]. This ecosystem of hardware and software technologies supports a growing interest in quantum computing and is evolving rapidly.
Despite its potential, integrating quantum computing tasks in practical systems is not without significant hurdles. Near-term practicality is held back by the variability in hardware resources and their supported software packages, lack of open-source frameworks that integrate classical and quantum workloads beyond python, latency between classical and quantum systems and sub-optimal job scheduling in shared-resource environments. Furthermore, automating workload execution on smaller-scale devices is challenging due to their experimental setting with specialized equipment and transient usage patterns. Unlike commercial platforms, such systems lack frameworks for abstractions, e.g., scheduling and execution of workloads while integrating with the hardware and software tools, particularly in HPC. Addressing these hardware and scheduling challenges is only part of the equation — an equally critical barrier lies in software tool integration. HPC benefits from optimization and parallelism provided by applications/frameworks, usually written in C/C++ (e.g. OpenMP [19] [20]).
To address the two critical issues of efficient scheduling and limited software tool integration, we introduce CONQURE, a novel fully featured modular quantum stack with HPC/Cloud integration. The contributions of this paper are as follows:
  • We create an open-source software framework for the management of quantum resources and workloads.
  • We evaluate our integrated software stack on simulated workloads.
  • We demonstrate functionality of CONQURE on an experimental ion-trap device.
  • We extend OpenMP to support quantum offloading (OpenMP-Q) and propose a design for reverse offloading within this framework.
[…]
II Background
The integration of quantum computing workloads into HPC requires interoperability between various resources, both classical and quantum, across the hardware and software stack. The fragmented landscape of frameworks supporting quantum resources, in particular, necessitates a middleware solution that can seamlessly integrate frameworks at different layers in the software stack, and specifically HPC software stacks while still providing Python compatibility. CONQURE’s architecture is designed to facilitate the automation of workloads on quantum devices today, eliminating concerns of software interoperability, development of supported QIR and hardware-specific toolchains.
II-1 Lack of Interoperability
Proprietary systems mandate the use of certain software libraries, which support specific hardware resources. While open-source solutions exist, they too come with assumptions about the hardware they are meant to be run on. While a number of well established frameworks exist and pipelines optimized for different kinds of devices have been developed, the average user is forced to make a decision which specific software library to use, based on the hardware they want to execute on. In contrast, our design allows the use of different hardware and software libraries via our custom translation layer.
II-2 Sequential Execution
Other middleware architectures, such as UQP [21], only support sequential operations as defined by their QIR. This limits their usability on near-term devices, which benefit from parallelization as a vital mechanism to mitigate the effect of qubit decoherence on overall noise. Also, the use of parallel gates is essential to get the most out of certain device architectures like neutral atom-based systems, which benefit greatly from the use of parallel operations on a large subset of qubits.
II-3 Lack of Software Tool Integration and Extension via Offloading Support with OpenMP
Quantum-assisted HPC provides novel opportunities to lower computational overhead for select algorithms. As such, traditional HPC can be complemented by quantum kernel to more efficiently solve highly complex real-world problems [22]. This introduces a middleware problem as software will be needed at the interface between classical HPC and low-level controlled quantum execution on QPU devices, i.e., connectivity and communication between classical and quantum devices is required. Past work introduced techniques like the distribution-aware Quantum-Classical-Quantum (QCQ) architecture that combines advanced quantum software frameworks with high-performance classical computing to improve quantum simulations [23] or Quantum-HPC Middleware [24]. However, most of the quantum programming languages and libraries are in Python (e.g., Qiskit).
CONQURE utilizes OpenMP, a powerful and widely-used framework for parallel programming, integral to general-purpose applications, ML, and HPC [25] with a high abstraction level to facilitate integration into classical computing. CONQURE provides novel support for quantum offloading via an OpenMP quantum extension, OpenMP-Q. OpenMP-Q provides target offloading to enable quantum computing as a QPU device option, similar to [26]. However, OpenMP-Q implements a pipe-based interaction between C++ and Python quantum libraries using LLVM-Clang [27] while also leveraging OpenMP’s multi-threaded model to execute multiple concurrent quantum tasks.
III Design
CONQURE is designed to provide an easy-to-deploy system that enables users to efficiently schedule workloads destined for quantum devices. This is done while also providing abstractions that simplify the submission of jobs and subsequent retrieval of results. While QisDAX [28] established a translation from Qiskit to DAX, CONQURE extends this by integrating a cloud-queuing mechanism and job persistence via an integrated database. CONQURE focuses on providing a robust infrastructure for handling job submissions, improving resource utilization, and ensuring scalability.
III-A CONQURE Stack
Figure 1: CONQURE Hardware-Software Stack
CONQURE employs a layered architecture designed for maximum flexibility across quantum hardware and software platforms. As shown in Fig 1, the system is designed with five modular layers that can be adapted for various architectures:
  • User Interface Layer: Provides programming frontends (Qiskit, Circ, Tket, etc.) for quantum kernel specification and classical-quantum workflow orchestration.
  • Translation Layer: Converts platform-agnostic quantum operations into hardware-specific instructions through interchangeable adapters. This layer abstracts vendor-specific compilation and optimization routines. This layer can be bypassed for use cases where the user has already created a workload with hardware specific routines.
  • Workload Manager: Combines cloud queuing, job scheduling, and resource management subsystems. Implements priority-based execution policies and hybrid workflow coordination between classical and quantum resources.
  • DB: Maintains job metadata, quantum circuit definitions, and execution results.
  • Quantum Control Layer: Device-agnostic interface for pulse-level control systems, designed to support diverse qubit technologies through pluggable drivers.
The architecture enables the replacement of components at any layer, i.e., users can substitute the quantum control layer to one designed for ion traps while designing circuits in Qiskit, or they can replace the scheduling subsystem without affecting higher-level APIs. This modularity ensures compatibility with emerging hardware technologies and evolving HPC software ecosystems.
III-B CONQURE Cloud Queue API
Figure 2: CONQURE Cloud Queue Architecture depicting API calls (red dashed box) and interactions between modules
III-B1 Compatibility with AWS Cloud Queue for Quantum Devices
CONQURE is designed as an interoperable alternative for AWS’ cloud queue for quantum devices [29], ensuring that researchers using AWS do not need to change their existing applications, but also providing missing components for an independent open-source stack for quantum researchers as opposed to commercial quantum service providers. The API calls for get_results() and create_work(), depicted in Listing LABEL:lst:sequre, are compliant with AWS’ Cloud Queue specification. Our API serves as a bridge to connect the translated user written code with the private cloud where the workload is run on either integrated QPUs or simulators. The red dashed box in Fig 2 signifies where the AWS-compliant interface CONQURE is situated.
However, in contrast to AWS, CONQURE is designed to provide services in a private cloud, i.e., an environment with authentication requirements that enables remote quantum kernel execution across research labs. Here, authentication needs to be maintained a priori to spawn CONQURE services.
III-B2 Scalability and Performance
Given that CONQURE is intended for use by researchers across labs, scalability was a key consideration in the API design. The system must be able to handle high volumes of requests without compromising performance. Fig 2 outlines the flow of data across different parts of the architecture. Our API supports asynchronous job submissions, allowing users to submit multiple jobs without needing to wait for their completion. This is particularly important here because job execution times are expected to vary significantly. SLURM [30] ensures efficient resource utilization with job queuing that can be tailored for each hardware device.
III-B3 Job Persistency and Tracking
It is important that the end-users are able to manage and track their workloads after submission. Persistent job storage allows users to retrieve information about submitted jobs, including workload information and target devices. Users can also query the status of jobs at any time to track their completion, or to retrieve historical results from prior executions.
III-B4 Modularity and Adaptability
CONQURE is designed as an intermediate layer between the software translation layers and the hardware control layer. It is crucial that it is able to integrate with different implementations of said software and hardware layers. Our API abstracts out this information and allows the user to send as a workload only those data items that are required by their downstream hardware control layer of choice.
$\begin{minted}$[
frame=lines, % Add a frame around the code
framesep=2mm, % Padding around the frame
breaklines=true, % Enable line breaking for long lines
% breakautoindent=true, % Optional: auto-indent wrapped lines
% breakindent=4ex, % Optional: set specific indent for wrapped lines
[…]
]{python}
from qiskit import QuantumCircuit, execute
from qiskit.providers.dax import DAX
import conqure
# GHZ Circuit definition in Qiskit
num_qubits = 4
ghz_circuit = QuantumCircuit(num_qubits)
ghz_circuit.h(0)
for i in range(num_qubits - 1):
ghz_circuit.cx(i, i + 1)
ghz_circuit.measure_all()
# QisDAX Translation Layer
dax = DAX.get_provider()
backend_name = dax_artiq_device
backend = dax.get_backend(backend_name)
backend.load_config("resources.toml")
dax_job = execute(ghz_circuit, backend,
shots=30)
workload = dax_job.get_dax()
# CONQURE UserClient
client = conqure.UserClient()
work_id = client.create_work(
workload=workload,
device_id=backend_name,
priority="LOW"
client.wait_until_done(work_id)
results = client.get_results(work_id)
# QisDAX Translation Layer
qiskit_result = dax_job.get_result_obj(results)
[…]
Support via OpenMP-Q}
Quantum programs consist of a sequence of gates, which are individual
[…]
execution of a quantum program typically involves repeated execution
of these sequences. When multiple quantum offload devices are
available, these gate sequences can be distributed across devices
based on partitioned data.
[…]
proposed quantum offloading strategy. The OpenMP-Q framework enables
reverse offloading, where the quantum kernel remains active while
classical computation runs asynchronously on the host. This reduces
task creation overhead and allows classical computations (e.g.,
parameter updates) to directly influence subsequent quantum gate
operations.
More specifically, results from one iteration of the quantum kernel
may trigger classical code execution (e.g., a solver), whose output is
relayed back to the quantum processor. These values are then
incorporated into the subsequent quantum gate executions, such as
phase-angle adjustments. This approach can be realized through
on-the-fly parametric pulse shaping within the FPGA that controls the
quantum device.
To facilitate quantum-host communication, we utilize a quantum class
[…]
enabling their communication with other MPI nodes during parallel
[…]
parallelized, while the quantum kernel continues execution with
updated angle values.
Furthermore, the classical solver could be implemented as a GPU kernel
or distributed across CPU cores on each node. The computed results are
aggregated to determine the best angle parameters before the next
quantum kernel iteration. MPI parallelization is optional; if MPI
calls are omitted, execution remains limited to a single
node. However, with MPI support, our model extends to multi-QPU
execution enabling result consolidation across quantum processors.
$\begin{lstlisting}$[caption={OpenMP-Q Single Quantum Offload},label=lst:qpu]
#pragma omp requires reverse-offload
void VQE(QuantumWrapper *c, int num_qubits, double angles[]) {
//add series of quantum gates, here: VQE
c->h(0); // Hadamard gate
for (int i = 0 ; i < num_qubits; i++)
c->ry(angles[i], i); // Y Rotation
for (int i = 0 ; i < num_qubits-1; i++)
c->cx(i, i+1); // CNOT gate
for (int i = 0 ; i < num_qubits; i++)
c->ry(angles[num_qubits+i], i); // Y Rotation
c->measure();
main(){
double angles[num_qubits] = init_angles(); // angles per VQE qubit
QuantumWrapper *c = new QuantumWrapper; // QuantumWrapper Class Object
for (i = 0; i < iterations ; i++) {
# pragma omp target device(Quantum) firstprivate(c) map(to: angles, qubits)
VQE(c, num_qubits, angles); // Quantum Gate Sequence
frequencies = c->execute();
# pragma omp target device (ancestor: 1) map (from: frequencies)
MPI_Broadcast(... frequencies ...); // distribute over nodes
angles = solve(frequencies); // classical
MPI_Allreduce(0 , ... angles ...);
pick_best_angles(angles);
Listing LABEL:lst:qpu_para demonstrates the execution of multiple parallel quantum tasks. By leveraging OpenMP directives, we schedule multiple quantum tasks according to the number of available OpenMP threads. The #pragma omp parallel directive enables multiple threads to execute distinct quantum circuits on separate QPUs or within quantum simulators. To ensure compatibility with OpenMP offloading semantics, each thread extracts its row from the global angles array into a temporary per-thread qpu_angles array, enabling thread-specific parameterization within the target region. Our language support generalizes to scenarios where multiple QPUs are available. For example, in parallelizing variational quantum algorithms with different Ansatzes [31], each node could be assigned a dedicated QPU with distinct initial angles supplied to evaluate multiple parametrized Ansatz strategies concurrently.
Listing 1: OpenMP-Q Multi Quantum Offload
double angles[num_qpus][num_qubits]= init_angles(); // angles per QPU and qubit: 2-D array
# pragma omp parallel
for(int i = 0 ; i < n_iterations ; i++) {
int qdev = omp_get_thread_num(); // one QPU per thread
double qpu_angles[num_qubits] = angles[qdev]; // angles per QPU
QuantumWrapper *c = new QuantumWrapper;
#pragma omp target device(Quantum) device_num(qdev) firstprivate(c) map(to: qpu_angles)
VQE(c, num_qubits, qpu_angles);
c->execute();
III-C Potential to Improve Iterative Classical-Quantum Algorithms
CONQURE’s modular, full custom architecture opens up opportunities to better optimize the pipeline for hybrid quantum-classical workloads, such as VQE and QAOA. While this paper focuses on the modularity, cloud integration and scheduling of CONQURE’s implementation, the design supports interactions between quantum and classical execution within the same job, which mitigates bottlenecks in hybrid workflows. (The implementation details are omitted due to space.)
III-C1 Challenges in Hybrid Iterative Workflows
Iterative algorithms such as VQE require several runs on a QPU with circuit parameters calculated using classical solvers. This can create substantial delays between the execution of quantum circuits. These idle periods significantly reduce hardware utilization efficiency. Data from Moses et. al [32] reveal that only a third of total operation time is used to run circuits on their Trapped-Ion System. The remaining overhead accounts for compiling circuits, retrapping of lost ions, etc.
CONQURE can prioritize those runs nearing convergence as these are highly sensitive to noise deviations in the device. Other optimizations specific to the hardware can also be exploited. For example, in ion-trap systems, those jobs with identical number of qubits can be prioritized to avoid the retrapping of ions between runs.
IV Implementation
IV-A CONQURE Cloud Queue
CONQURE integrates components from several existing frameworks:
  • User Code: Qiskit provides high-level abstractions for building quantum circuits.
  • Translation Layer: QisDAX bridges Qiskit with DAX, allowing execution on ion-trap devices.
  • Quantum Control Interface: ARTIQ facilitates low-level control of FPGAs for ion-trap hardware.
  • SLURM: Enables job scheduling and queuing.
  • Flask: Manages the REST API for communication between clients and servers.
  • MariaDB: Stores workload related information like job status, results, etc.
It should be stressed again that CONQURE is built in a modular fashion such that its components can easily be replaced for different hardware or software environments. Our current implementation uses the above mentioned frameworks. The typical workflow is as follows:
  • The user must first generate a workload and specify the target device. In Listing LABEL:lst:sequre, this workload is generated by QisDAX through its get_dax() method.
  • The user invokes the CONQURE UserClient class’ create_work to create a job on the server by sending this workload as well as information about the target device as well as a job priority.
  • On the master server, this workload is pushed into a new entry in the database and a unique job_id is returned to the user. This ID they can then be used to track the job status and retrieve results once the job has executed.
  • Once the job has completed, the results are pushed into the central DB, and the completion status of the job is logged in the entry.
  • The user invokes the CONQURE UserClient class’ get_results to retrieve the raw data from the DB. This is the data given by the quantum control layer and may need to be parsed to be usable. In the implementation shown in Listing LABEL:lst:sequre, the data retrieved matches ARTIQ’s output requirements. It is subsequently transformed into a Qiskit result object to ensure compatibility with any downstream Qiskit code.
Figure 3: CONQURE: OpenMP Quantum Offloading Pipeline (OpenMP-Q)
Listing 2: OpenMP-Q Quantum Offload Example
std::vector<std::vector<double>> angles = std::vector<std::vector<double>> (N, init_angles());
#pragma omp parallel
qdev = omp_get_thread_num();
for(int i = 0 ; i < K ; i++) {
device_angles = angles[qdev];
QuantumCircuitWrapper *c = new QuantumCWrapper(qubits);
# pragma omp target device(Quantum) firstprivate(c) map(tofrom : device_angles [0:num_qubits])
VQE(c, num_qubits, device_angles);
frequencies = c->execute();
angles[qdev] = solver(frequencies);
Listing 3: OpenMP-Q, LLVM OpenMP Offload Updates
targetKernel(Int64_t DeviceId, KernelArgsTy *KernelArgs) {
if(DeviceId == Quantum) {
// (1) Retrieve quantum circuit object pointer
c = (QuantumCircuitWrapper*) KernelArgs->ArgBasePtrs[n]; // n = 0 (serial) or 1 (parallel)
for (int32_t I = 0; I < KernelArgs->NumArgs; ++I){
if (KernelArgs->ArgTypes[I] & OMP_TGT_MAPTYPE_TO) {
// (2) Parse mapped device angles from ‘map(to: device_angles)‘ and serialize
processDataMapTo(KernelArgs->ArgBasePtrs[I])
// (3) Execute user-defined offload region -- populate quantum gate sequence
target(...)
if (DeviceId == Quantum) {
// (4) Generate and execute Python script
c->execute_python();
// (5) Write back result from python to device_angles via ‘map(from: device_angles)‘
processDataMapFrom(KernelArgs->ArgBasePtrs[I]))
IV-B OpenMP-Q
CONQURE contributes a common embedding library with an integration into LLVM frontends [27], which connects to quantum intermediate representations (QIR) for quantum gates via our OpenMP-Q extension. Figure 3 shows the complete pipeline of the C++/python bindings with OpenMP in CONQURE as follows.
  • Consider Listing LABEL:lst:qpu_2. The user writes a simple OpenMP program to execute quantum tasks on multiple QPUs using the parallel directive. The target directive is used to specify the sequence of quantum gates to execute. The QuantumWrapper class is added to the OpenMP shared library allowing users to create a pointer to the quantum object and pass it to the OpenMP runtime using the firstprivate clause (Fig 3-1).
  • At runtime, the device clause is extended in OpenMP-Q to include a quantum device ID identifying the offloaded target as a quantum device. Listing LABEL:lst:llvm demonstrates the required updates within OpenMP runtime offload interface to integrate OpenMP-Q. #pragma omp target makes a targetKernel function call with kernel argument parameters. From the numbered points marked as comments in the listing LABEL:lst:llvm we further elaborate the updates within the targetKernel function:
    The pointer to the user initialized quantum object inside the firstprivate clause is extracted using the kernel arguments.
    The mapped data to the quantum device is parsed, serialized and stored within the object (to be passed as system arguments to the Python script if needed).
    The user-provided offload code is executed (see Fig. 3-2). This offloading code defines the sequencing of quantum operations within the quantum object, enabling the generation of a Python script that specifies a quantum task. Thread IDs at OpenMP runtime are used to map different tasks to individual QPUs and mapping relevant data.
    The OpenMP offload interface immediately generates and executes this script. The implementation uses a Operating System (OS) pipe between processes to run the CONQURE task alongside the quantum object data (Fig 3-2 & Fig 3-3).
    Results from the executed Python script are de-serialized, parsed and updated into the values mapped back to host.
  • Nested target regions, in case of reverse offload (see Sect. 2), would utilize the bidirectional communication functionality of Unix pipes (Fig 3-4). (Notice: Due to incompatibility of reverse offload functionality with the current LLVM OpenMP versions, this feature is a design for future implementation subject to OpenMP extensions within LLVM in the first place. We include it here to discuss the benefits of such reverse offloads for quantum.)
  • The OpenMP-Q clause can also be extended to add a Python script instead of a gate sequence on a quantum circuit. In this case, OpenMP-Q executes the script directly, passing the mapped data in the same way, i.e., as a serialized string, to the Python executable.
Listing LABEL:lst:py presents a summary of the Python script generated by OpenMP-Q for the code in Listing LABEL:lst:qpu_2. OpenMP-Q offers users a flexible interface to insert quantum gate sequences in any desired form.
Listing LABEL:lst:json illustrates a standard Python-to-C++ message that conveys the frequencies of observed classical qubit states. This message is parsed at runtime by OpenMP, storing the evaluated frequencies in an array. These values are then used to compute updated angles via the angle solver (see Listing LABEL:lst:qpu_2). Unobserved states are assigned a frequency of 0.
Listing 4: Generated Python Script: Circuit Execution
if __name__ == "__main__":
circuit = QuantumCircuit(4)
circuit.ry(2.858849, 0)
circuit.ry(1.445133, 1)
circuit.ry(2.136283, 2)
circuit.ry(2.293363, 3)
circuit.cx(0, 1)
circuit.cx(1, 2)
circuit.cx(2, 3)
circuit.ry(1.445133, 1)
circuit.ry(2.136283, 2)
circuit.ry(2.293363, 3)
circuit.ry(1.043242, 3)
circuit.measure_all()
backend = Aer.get_backend(’statevector_simulator’)
counts = execute(circuit, backend, shots=100).result().get_counts()
counts = json.dumps(counts)
print(counts)
Listing 5: Frequency of different qubit states (4 qubits)
{"1011": 8, "0011": 7, "0111": 14, "1001": 7, "0101": 3, "1110": 1, "1111": 60}
V Results
The CONQURE framework was deployed and tested using simulators and an ion-trap quantum device at the Duke Quantum Center [33], which we have access to. The following sections detail the experiments performed and results obtained.
V-A CONQURE Cloud Queue
We evaluate the two key methods within the UserClient class that manage the creation of jobs at the backend by sending a workload and subsequent retrieval of results from the database. We experimented with different sized GHZ state preparation circuits [34] to estimate the overhead. GHZ State preparation circuits were used here as their circuit size increases linearly with qubit count. All experiments were repeated 1,000 times. These tests reflect the data collected on a local instance of CONQURE with a simulator, i.e. the user’s code is run on the same machine that the CONQURE server is hosted on. Latency was measured on a system with a configuration as indicated in Listing LABEL:lst:spec with Python version 3.9.18 and LLVM version 20.0.0 enhanced by and recompiled for our OpenMP-Q extension.
$\begin{minted}$[
frame=lines,
framesep=2mm,
breakafter=:, % Suggest breaking *after* the colon if possible
breaklines=true, % *** Enable automatic line breaking ***
breakautoindent=true, % Indent wrapped lines automatically based on context
breakindent=2ex, % Amount of extra indentation for wrapped lines (adjust as needed)
]{yaml}
OS : Ubuntu 22.04.5 LTS
CPU : Intel(R) Core(TM) i9-9900 CPU @ 3.10GHz (8 cores)
GPU : NVIDIA GeForce RTX 2080 Ti
RAM : 16GB
Swap: 4GB
Disk: 1TB
[…]
(x-axis) and backend configurations (legend) indicating latency in ms
(y-axis). We observe an average response time of 12.69ms and 12.82ms
when targeting a real device and a simulator, respectively. There are
small deviations given by box plots indicating median, quartiles and
minimum/maximum measurements obtained. Notice that this API call
triggers the execution, i.e., it mainly registers the activity in the
database and spawns a corresponding asynchronous job.
$\begin{figure}[h]$
$\begin{center}$
[…]
(x-axis). This metric is the product of number of shots and number of
qubits measured as both factors affect latency. Our results show a
[…]
is to be expected as the sheer amount of data being transferred is
lower given that results are simply read out from the database after
job completion.
$\begin{figure}[h]$
$\begin{center}$
[…]
experimental ion-trap device at Duke Quantum Center with 6 qubits. At
the time of testing, this system only supports single qubit
operations. To simulate the execution of these circuits to gauge
overhead and latencies, we replace any two qubit operations by a
sequence of single qubit operations, on both qubits, equal in time to
the the two qubit operation. More specifically, CNOT gates were
replaced by an RY and an RX gate on the target qubit, followed by two
Hadamard and two NOT gates on both qubits, followed by another RY and
RX on the target. This changes the unitary of the circuit and, hence,
[…]
from this modified unitary.
[…]
Next, we assess the efficacy of our OpenMP-Q extension
through a VQE experiment. VQE involves estimating the
minimum eigenvalue of a Hamiltonian by optimizing the parameters of an
ansatz circuit. The expectation value of the Hamiltonian is calculated
by measuring the circuit and evaluating over a set of simpler terms,
such as Pauli strings, as a first approximation. The parameters of the
ansatz circuit is then optimized using a classical solver for the next
iteration, and this process is repeated until convergence. However,
this approach faces two problems, that of getting trapped into local
minima and slow convergence due to barren plateaus. To mitigate these,
multiple initial states are chosen, and the lowest measured
expectation value is returned. This process is typically done
sequentially.
$\begin{figure}[h]$
$\begin{center}$
[…]
We test our system with a parallel VQE implementation, similar to the
[…]
the convergence of various QPU runs with randomized initial
parameters, i.e., each colored curve corresponds to a different set of
ansatz parameters resulting in decreasing cost (y-axis). We schedule
these VQE runs (a) serially and (b) in parallel on simulators, in the
latter case to show the potential of parallelization over QPUs as a
concept to optimize convergence time, i.e., the different runs
themselves are parallelized.
We analyze the runtimes across different number of runs for the same
VQE problem, both with and without our Multi-Q Offloading
[…]
to 6 VQE runs (x-axis) serially (red) and OpenMP-Q parallelized
(blue). Experimenters were repeated 200 times and showed minimal
variations (indicated by barely visible whiskers for 1st and 3rd
quartiles). We observe significant speedups as the number of threads
is increased. Given that the number of available QPUs is equal to the
number of threads, a single VQE run using our OpenMP-Q Offload
standard takes 38sec. But serially executing 6 VQE kernels takes
228sec, whereas running 6 runs in parallel requires only 71sec. This
[…]
$\begin{figure}[h]$
$\begin{center}$
[…]
VI Related Work
Research into the integration of quantum computing into HPC environments is a critical endeavor, driven by the objective to realize quantum advantage on computationally complex kernels in classical terms, which can be solved more efficiently on quantum devices.
Saurabh et al. proposed a conceptual middleware solution [24] to build a foundation for the future of HPC middleware systems. Mantha et al. built upon this foundation with a middleware solution designed to manage classical and quantum resources at the application level [36]. CONQURE’s modularity and private cloud complement this design while focusing on providing a comprehensive framework for managing quantum and hybrid classical-quantum workloads across hardware platforms.
UQP [21] introduced a platform designed for the integration of HPC environments with quantum accelerators. It contributed a novel ISA understood by UQP, which is translated from a Quantum Intermediate Representation (QIR). CONQURE also aims to integrate quantum computers into HPC environments but focuses on a practical near-term implementation that integrates frameworks used in HPC (OpenMP) with a private cloud.
The scheduling and management of resources and workloads is critical in the practical deployment of these tools. Qoncord [37] tackled the problem of workload scheduling in cloud environments and proposed a novel scheduling framework that capitalizes on the differences in approximation errors over in different phases of VQA. It splits VQA runs into distinct exploratory and fine-tuning phases, and identified that higher noise was acceptable in the exploratory phase and exploited this to schedule it with lower priority. Unlike Qoncord’s focus on noise resilience, CONQURE focuses on scalability, job persistence and tracking with an emphasis on hybrid classical-quantum workloads.
CONQURE also extends quantum task execution to HPC by implementing OpenMP-Q, a quantum extension to OpenMP. Historically, HPC has continually evolved by embracing new processing paradigms and by successfully integrating special-purpose accelerators to enhance performance. In a similar vein, incorporating quantum accelerators into HPC workflows presents a promising path forward for tackling problems beyond the reach of classical systems. [38] discusses quantum integration strategies to build a simplified CPU-QPU (Quantum Processing Units) execution model to integrate into current and future HPC system architectures.
Among the many parallel programming models used in HPC, MPI and OpenMP stand out as the most widely adopted frameworks for distributed and shared-memory parallelism, respectively. By building on OpenMP with its existing support for accelerators such as GPUs for computational kernels, CONQURE provides a natural and portable path for extending existing HPC applications to leverage quantum acceleration with minimal disruption to existing codebases. [26] presents the closest related effort in terms of the OpenMP-Q contribution in CONQURE, to the best of our knowledge. Their work extends OpenMP to support quantum offloading through function calls that create and measure quantum registers and apply a fixed set of single- and two-qubit gates. These circuits are then transpiled into QASM or QIR for execution. While their approach addresses a similar problem domain, not results are reported for this poster. Their work also lacks the generality and scalability required for broader quantum-classical integration into the software stack that CONQURE provides with its full LLVM implementation.
Furthermore, CONQURE introduces OpenMP-Q, which dynamically generates Python scripts at runtime, enabling interoperability with a wide range of external quantum frameworks, not just fixed simulators. Additionally, OpenMP-Q supports bidirectional data exchange between the host and quantum code through a shared quantum object, enabling runtime feedback and classical reuse of quantum results. This design is especially valuable for hybrid quantum-classical workflows such as VQE and QAOA, where iterative refinement based on quantum output is essential.
VII Conclusion
The integration of quantum hardware and software ecosystem presents significant challenges due to its fragmented nature and lack of interoperability. Similar challenges are faced when trying to integrate QPUs into HPC workflows, coupled with the lack of standardized interfaces. In this work, we introduced CONQURE, a novel, open-source co-execution framework that fill these gaps.
CONQURE is designed as a modular, five-layer framework that includes user interfaces, a translation layer, workload management, job scheduling, database persistence and a quantum control layer, enabling compatibility across diverse software and hardware platforms, which was tested in simulation and on an ion trap device. We also introduced OpenMP-Q, an addition to the OpenMP standard, which enables seamless offloading of quantum kernels onto QPUs with minimal integration effort for the user.
[…]
Acknowledgment
This work was supported in part by NSF awards MPS-2410675, PHY-1818914, PHY-2325080, OMA-2120757, and CISE-2316201. We would like to thank Andrew Van Horn from Duke Quantum Center for his support in the testing of CONQURE on real devices.